OpenStudy (anonymous):

7th grade math help @cj49

OpenStudy (anonymous):

Chris’s hike started at an elevation of –12 m. His hike finished at an elevation of 235 m. How much did Chris’s elevation change?

OpenStudy (anonymous):

@cj49

OpenStudy (anonymous):

@HassanAliPK

OpenStudy (anonymous):

I need help with 3 more problems after this please there is my problem:)

OpenStudy (anonymous):

The change in elevation is the final elevation minus the initial elevation. In this case: 235--12 (two minus make a plus) 235 + 12 = 247m

OpenStudy (anonymous):

Ravi had $11 in his piggy bank on Monday. He deposited $1.30 into the piggy bank on Tuesday, and removed $2.85 on Wednesday. How much was left in Ravi’s piggy bank?

OpenStudy (anonymous):

deposit means you add to the money in the bank (+) remove means you minus money in the bank (-) Ravi starts with $11 plus $1.30 minus $2.85 11 + 1.30 - 2.85 = $9.45

OpenStudy (anonymous):

This morning, Robyn wrote a check in the amount of $35.80. This brought her balance down to –$18.25. Suppose n represents Robyn's balance before she wrote the check. Which equation can be used to represent this situation? A. n – 18.25 = 35.80 B. n – 35.80 = –18.25 C. n + 35.80 = –18.25 D. n – 18.25 = –35.80

OpenStudy (anonymous):

B. n – 35.80 = –18.25 This is because Robyn removed $35.80 from here initial bank balance n. This left her with a negative balance of $18.25

OpenStudy (anonymous):

which letter is it???

OpenStudy (anonymous):

I put the letter at the beginning of my answer...

OpenStudy (anonymous):

A zoologist is recording the loss of wolves in a her state. She notes the number of wolves, w, in the state on January 1. One year later, there were 84 wolves in the state, which is 23 fewer wolves than were in the state a year earlier. Which equation represents this situation? A. w + 23 = 84 B. w – 23 = 84 C. 23 – w = 84 D. w – 84 = –23

OpenStudy (anonymous):

I think you should try to work this one out.

OpenStudy (anonymous):

can u at least help me

OpenStudy (anonymous):

is it B.

OpenStudy (anonymous):

What makes you think it is B? Can I see your steps?

OpenStudy (anonymous):

Yes, it's (B)

OpenStudy (anonymous):

Ok but we are here to help you learn. It is ok if you're not sure how to start the problem. When the problem states 23 fewer wolves, do you think that means you should minus 23 or add 23?

OpenStudy (anonymous):

Is my answer right???

OpenStudy (anonymous):

d is right